# A Comprehensive Evaluation of Semantic Relation Knowledge of Pretrained Language Models and Humans

### v1

@@ Abstract
Recently, much work has concerned itself with the enigma of what exactly PLMs (pretrained language models) learn about different aspects of language, and how they learn it. One stream of this type of research investigates the knowledge that PLMs have about semantic relations. However, many aspects of semantic relations were left unexplored. Only one relation was considered, namely hypernymy. Furthermore, previous work did not measure humans’ performance on the same task as that solved by the PLMs. This means that at this point in time, there is only an incomplete view of models’ semantic relation knowledge. To address this gap, we introduce a comprehensive evaluation framework covering five relations beyond hypernymy, namely hyponymy, holonymy, meronymy, antonymy, and synonymy. We use six metrics (two newly introduced here) for recently untreated aspects of semantic relation knowledge, namely soundness, completeness, symmetry, asymmetry, prototypicality, and distinguishability and fairly compare humans and models on the same task. Our extensive experiments involve 16 PLMs, eight masked and eight causal language models. Up to now only masked language models had been tested although causal and masked language models treat context differently. Our results reveal a significant knowledge gap between humans and models for almost all semantic relations. Antonymy is the outlier relation where all models perform reasonably well. In general, masked language models perform significantly better than causal language models. Nonetheless, both masked and causal language models are likely to confuse non-antonymy relations with antonymy.
@@ 1 Introduction
The present study tries to establish to what extent they are able to do so. We extend the existing methodology by introducing a new evaluation framework. We cover six relations, namely hypernymy, hyponymy, holonymy, meronymy, antonymy, and synonymy. Our framework is also the first to shine a light on previously understudied properties of semantic relations and meta-relations. Two kinds of comparisons are considered. First, we compare models against humans on the same task, so that we can quantify the difference with the theoretically achievable ceiling. Second, we compare two families of models, which differ in the pretraining tasks used. Within families, we consider different sizes ranging from millions to billions in parameters. These comparisons allow us to identify which factors facilitate the acquisition of semantic relation knowledge by the models. In sum, our evaluation framework deeply explores both well-studied and previously unexplored properties, and it uses new metrics in a new comparative experimental setting. By doing so, it adds both depth and width to the current knowledge about the quality of semantic relation knowledge that can be acquired by today’s PLMs.
@@ 2.1 Probing for Hypernymy
In contrast, Ettinger [16] pioneered the study of hypernymy with prompt-based probing. In prompt-based probing, responses from a model are elicited using a prompt, a textual string with slots, all of which are unfilled. An example prompt for hypernymy is “a [w] is a [v] where [w] is a slot for the word given (called the target word here), whereas [v] will be predicted by a model. As the second word, which is called relatum, depends on the target word and the relation described. A probe is a prompt where the target word has been filled in. For example, Ettinger used the string “a robin is a [v] as a probe. If a PLM has learned hypernymy well, it should be able to predict hypernyms of “robin” for [v], such as “bird”.
@@
Hanna and Mareček [21] further developed Ettinger’s and Ravichander et al’s prediction task, exploiting the fact that a semantic relation can be expressed by multiple prompts. For the prompts, Hanna and Mareček adopted some lexico-syntactic patterns known from previous work  [22] to be effective at retrieving word pairs in hypernymy and hyponymy. Such prompts include “my favorite [w] is a [v], “a [w], such as a [v], “a [w] is a type of [v], “a [w] is a [v]. Hanna and Mareček dealt with the problem of ambiguous target words in a different way from Ravichander et al [50]. They attached an example sentence from SemCor [30] to each probe, whether the target word was ambiguous or not. The example sentence was chosen in such a way that the target word was in a specific sense, namely the first WordNet sense whose hypernyms include the gold hypernym. In Hanna and Mareček’s experiments, BERT reached an accuracy of 0.48 in the best setting. Contrary to intuition, the accuracy dropped by about 0.05 when the example sentences were used.
Another modification of the original task is the evaluation data Hanna and Mareček used, which came from category norms. Category norms are items that humans judge to be subtypes of a category given to them. For instance, one category norm for “fish” is “tuna”; “trout” and “salmon” are others. Cohen et al [11] were the first to collect category norms. In their experiment, each participant wrote down category norms included in one of 43 categories given to them, as many as came to mind in 30 seconds. Cohen et al’s original category norm collection was later expanded by Battig and Montague [3], who added 13 new categories (for a total of 56 categories and 2,082 category-category norm pairs). Another colletion effort by Overschelde et al [46] added 14 more new categories (for a total of 70 categories and 1,983 category-category norm pairs).
@@
Hanna and Mareček applied the hyperynym-based part of Battig and Montague’s category norm data to their probing experiments in the obvious way, using category norms as target words and categories as gold relata. They discarded categories and category norms that were tokenized into multiple tokens, ending up with a total of 863 norms paired with 25 categories.
In summary, based on the above studies we can conclude that BERT is able to predict hypernyms of target words to a reasonable degree. However, several interesting avenues have been left unresearched. The most obvious of these is that there are several other established semantic relations apart from hypernymy.
@@ 2.2 Relations beyond Hypernymy
Hypernymy is only one type of semantic relation, although it is an essential one. Other semantic relations also have been a long-standing research topic in psychological, theoretical, and computational linguistics. Studies revealed that the distinction between semantic relations is non-trivial for both humans and models [9, 10, 27, 56, 44, 2, 65].
In psycholinguistics, Chaffin and Clark [9] researched the similarities and difference between several semantic relations, as perceived by humans. The experiment used a semantic sorting task, where participants are instructed to group together 31 word pairs, each representing a particular semantic relation. The relations came from five broad categories: contrast (including antonymy), similars (including synonymy), class inclusion (including hypernymy), part-whole (holonymy), and case relations (such as the agent-instrument relation and the agent-action relations, exemplified by “farmer”/“tractor” and “dog”/“bark”). The results showed that the human subjects were able to distinguish contrast (including antonymy) most easily from the other four relations. Similars (including synonymy) and class inclusion (including hypernymy) formed a second cluster, whereas case relations and part-whole formed a separate cluster each.
The similarities between hypernymy and holonymy have been extensively discussed in the semantic literature [12, 64, 27]. Both Cruse and Winston et al pointed out that hypernymy and holonymy are similar in that they both involve division and inclusion. Word pairs related by hypernymy and holonymy always consist of a word refering to an entity that undergoes division, with the other word referring to the result of that devision, whether as a part or a subclass. Joosten considered both relations under the term denotational inclusion. The similarity between the relations becomes even more obvious when collective nouns are involved. For example, we can say that a table is a kind of furniture, and we can also say that it is a part of furniture, expressions typically associated with hypernymy and holonymy [27].
@@
Synonymy and hypernymy are also closely related, as was confirmed during the creation of Hyperlex [62]. Hyperlex is a lexical resource of semantic relations (hypernymy, hyponymy, meronymy, synonymy, antonymy, co-hypernymy [^2] Co-hypernymy is the relation between two concepts that share a hypernym, such as “hawk” and “robin”.), holding between 2,616 word pairs of nouns or verbs333290 unrelated word pairs also exist.. The original pairs in Hyperlex were sampled from WordNet [39] and the University of Southern Florida Norms dataset (USF) [43]. Three human checkers were asked to verify whether the semantic relation holds for each pair sampled; the pair was discarded unless two of them agreed that it did. Next, different crowd workers were asked to assign a score to the remaining pairs, indicating the degree to which the pair satisfies hypernymy. Note that for those pairs that were related in a non-hypernymy relation, the crowd workers should assign a low score. However, the humans’ score for synonymy was close to that for hypernymy pairs if the two words were close to each other in the WordNet hierarchy (i.e., separated by at most two levels).
2.3 Symmetry, Asymmetry and Prototypicality
When assessing a model’s knowledge about semantic relations holistically, it is necessary to consider not only if the model uses the semantic relations correctly, but also to what extent it learns which specific properties the semantic relations have.
Symmetry and Asymmetry
One such property is symmetry, which is defined as follows: If a word pair $(w_{1},w_{2})$ is in a symmetric relation, then the reverse pair $(w_{2},w_{1})$ is also in the relation. A related property is asymmetry, the opposite of symmetry: if the word pair $(w_{1},w_{2})$ is in an asymmetric relation, then the reverse pair $(w_{2},w_{1})$ is not in the relation. It has been experimentally shown that modeling asymmetry improves the performance of semantic relation classification tasks [19].
Symmetry and asymmery are also properties of some factual knowledge relations. For instance, the factual knowledge relation “is a sibling of” is symmetric. Mruthyunjaya et al [42] proposed metrics in order to assess whether models learn properties of such factual relations, including symmetry. They used prompt-based probing and defined the concept of reciprocal elicitation: for any word pair (target word $w$ and relatum $v$) that forms a symmetric relation, the model should respond with the relatum, when given the target word, and also respond with the target word, when given the relatum. For the probe “Bart Simpson is a sibling of [v], they expected models to predict “Lisa Simpson”, and for the converse probe “Lisa Simpson is a sibling of [v], to predict “Bart Simpson” .
For every such word pair $(w,v)$, a prompt $p$, and a model $m$’s top $k$ items in the response $m_{k}(w,p)$, given $w$ and $p$, the symmetry score is the average of $\mu_{k}(w,v,p)$ over all $(w,v)$ and $p$.
@@ 2.3 Symmetry, Asymmetry and Prototypicality
where $\mathbb{I}[P]$ is the indicator function that becomes one when $P$ is true and zero otherwise. In contrast, for asymmetric relations there should be no reciprocal elicitation but only forward elicitation: the model should only respond with the relatum given the target word, not with the target word given the relatum. The asymmetry score is the average of $\alpha(w,v,p)$ over all $(w,v)$ and $p$.
$\alpha_{k}(w,v,p;m)=\mathbb{I}[v\in m_{k}(w,p)]\times\mathbb{I}[w\notin m_{k}(v,p)]$ (2)
Both scores range between zero and one. Intuitively, the symmetry score can be interpreted as the probability of reciprocal elicitation for a semantic relation $r$. The asymmetry score, on the other hand, can be interpreted as the probability of only forward elicitation happening for $r$. For symmetric relations, a high symmetry score means that agents were able to detect symmetry; for asymmetric relations, a high asymmetry score similarly means that agents were able to recognize the asymmetric nature of the relation.
Mruthyunjaya et al’s results showed that, for symmetry, BERT outperforms even GPT-3 [6]. For asymmetry, GPT-3 outperforms BERT.
@@
There have been theoretical discussions of the prototypicality of holonymy. Taylor [59] and Tversky [61] presented top-down accounts, which emphasize that the whole (holonym) is intrinsic in the conceptualization of the part (meronym). They therefore predict that the relationship between the whole and its mandatory parts should be tighter than between the whole and its optional parts. A building is a structure with walls, and walls are defined by their function within a building. This makes “building” a prototypical holonym of “wall”. But not all holonymy pairs are well described by these accounts, because some optional parts also play an important role. This is acknowledged in the bottom-up accounts [31, 38, both cited by Joosten [27]], who state that a whole is formed by assembling a number of other individual wholes, each of which has a separate existence outside the holonymy relation. For example, “sky” is a typical holonym of “cloud” (indeed, in our forthcoming experiments, it happens to be the holonym most frequently mentioned by humans). However, in sunny weather, the sky can be cloudless, so clouds are not intrinsic to the sky. This seems to make bottom-up theories descriptively more adequate, although neither proposes a prototype prediction mechanism.
@@ 3 Methodological Considerations
The majority of the previous research studied hypernymy. Beyond hypernymy, there is also much theoretical and experimental knowledge about semantic relations. Additionally, the literature has established several facts about meta-relations. However, when it comes to practical investigations of model behaviour, it is always only hypernymy that has been studied, even though it is merely one semantic relation amongst many.
Additionally, we have seen that the relevant theoretical literature has extensively studied meta-relations such as confusability between semantic relations (cf. section 2.2). In stark contrast with this, the methodology previously used is unable to establish the degree to which a model mistakes one semantic relation for another, and methodologies for other meta-relations are non-existent. This leaves us with an incomplete understanding of the nature of semantic relations, and of the knowledge that PLMs have about semantic relations.
Therefore, we design new metrics for measuring prototypicality (a property of relations) and distinguishability (a meta-relation). Using these new metrics, and the established ones for symmetry and asymmetry, we study hyponymy, holonymy, meronymy, antonymy and synonymy, as well as hypernymy. We further provide a direct comparison between models and humans on the same task. Such a human ceiling will allow us to interpret the performance of models more meaningfully.
@@
Another recurrent problem for all probing experiments is that most target words from any source, are naturally sense-ambiguous words. Previous experimental attempts to deal with sense ambiguity are suboptimal. Ravichander et al limited the target words they use to the unambiguous words. This limits the research focus to an artificial subset of all possible words and relations, and has the practical disadvantage of considerably reducing the number of target words one can use. The other existing solution is to provide context of the target word sense, for example in the form of example sentences, as Hanna and Mareček did, but this empirically harmed the performance of models. Our methodology offers an alternative solution to this problem.
@@
Some previous experiments on relatum prediction assumed that there is only one correct relatum for each tuple. Hanna and Mareček [21] acknowledge that “orange” has two gold relata: “color” and “fruit”. They therefore defined two separate gold tuples for this probe: (“orange”, HYP, “color”) and (“orange”, HYP, “fruit”). However, when calculating the accuracy scores, they consider only the first item in the responses for both tuples, and then average over the two tuples. In this setting, it is theoretically impossible for a model to achieve the full score (1) for any sense-ambiguous probe, even if the model had the ability to predict both relata. Whether the model predicts [“color”, “fruit”] or [“fruit”, “color”], the accuracy score is always 0.5555In the general case, the highest achievable accuracy score for target words with $n$ relata is $\frac{1}{n}$..
@@
A confounder in the interpretation of numerical results is the use of definite and indefinite determiners in the probes. Previous researchers routinely used probes that include indefinite determiners, such as Ettinger’s (2020) “a robin is a [v]. The English indefinite determiner changes its form from “a” to “an” if the following word’s pronunciation starts with a vowel. Choosing “a” or “an” in a probe before the [v] slot would therefore bias the prediction by models towards relata with an initial vowel or consonant.
Ettinger ran experiments using both types of determiners, comparing pairs of probes differing only in the determiners used. Manual inspection of the responses showed that BERT indeed always adhered to the morphophonetic rule. For example, given the probe “a hammer is a [v], BERT-large predicts [“hammer”, “tool”, “weapon”, “nail”, “device”], whereas given the probe “a hammer is an [v], it predicts [“object”, “instrument”, “axe”, “implement”, “explosive”]. This suggests that the BERT models were able to utilize the grammatical information contained in the determiner as a clue.
Ravichander et al, when faced the problem of which indefinite determiner to place before [v], chose to always use the determiner that morphophonetically fits with the gold standard answer. For example, the determiner in the probe “a moth is an [v] was chosen to be “an” exactly because the gold standard answer was “insect”. However, when the determiner acts as a clue for the model, it becomes impossible to disentangle how much of the results is due to the model’s semantic knowledge and how much to the clue. Choosing randomly also doesn’t solve the problem. Our solution uses both probes with “a” and probes with “an” and merges the results in a statistical manner.
@@
Another understudied aspect is the comparison between different types of PLM models on the semantic probing task. BERT, the only model studied so far, is a masked language model (MLM). MLMs are pretrained on the masked language modeling task, in which a model is asked to recover tokens in a given sentence that are randomly masked. But recently, causal language models (CLMs) such as OPT [66] and Llama [60] have shown high performance in many tasks and thus gained attention. CLMs are pretrained on next-token prediction, the task of predicting the rightmost word given a sequence of words. The difference in pretraining tasks means that the models make their decisions based on different kinds of information. MLMs consider the context of both sides of the masked word, while CLMs consider only the preceding context. Previous studies in a number of tasks found a large influence of the type of context used in PLM pretraining on performance. For instance, for factual knowledge, MLMs have been found to be superior over CLMs [49, 8, 42]. We are the first to study the role of bidirectional contexts in the recognition of semantic relations. To be fair to both MLMs and CLMs, we have designed all our prompts in such a way that they end with the slot [v].
@@
We always first establish human performance for each task and then use it as the measuring stick for models’ performance [^6] Except in the case of asymmetry, as we will explain in what follows..

### v4

@@ Abstract
Recently, much work has concerned itself with the enigma of what exactly pretrained language models (PLMs) learn about different aspects of language, and how they learn it. One stream of this type of research investigates the knowledge that PLMs have about semantic relations. However, many aspects of semantic relations were left unexplored. Generally, only one relation has been considered, namely hypernymy. Furthermore, previous work did not measure humans’ performance on the same task as that performed by the PLMs. This means that at this point in time, there is only an incomplete view of the extent of these models’ semantic relation knowledge. To address this gap, we introduce a comprehensive evaluation framework covering five relations beyond hypernymy, namely hyponymy, holonymy, meronymy, antonymy, and synonymy. We use five metrics (two newly introduced here) for recently untreated aspects of semantic relation knowledge, namely soundness, completeness, symmetry, prototypicality, and distinguishability. Using these, we can fairly compare humans and models on the same task. Our extensive experiments involve six PLMs, four masked and two causal language models. The results reveal a significant knowledge gap between humans and models for all semantic relations. In general, causal language models, despite their wide use, do not always perform significantly better than masked language models. Antonymy is the outlier relation where all models perform reasonably well.
@@ 1 Introduction
The present study tries to establish to what extent they are able to do so. We extend the existing methodology by introducing a new evaluation framework. We cover six relations, namely hypernymy, hyponymy, holonymy, meronymy, antonymy, and synonymy. Our framework is also the first to shine a light on previously understudied properties of semantic relations and meta-relations. Two kinds of comparisons are considered. First, we compare models against humans on the same task, so that we can quantify the difference with the theoretically achievable ceiling. Second, we compare two families of models, which differ in the pretraining tasks used. Within families, we consider different sizes. These comparisons allow us to identify which factors facilitate the acquisition of semantic relation knowledge by the models. In sum, our evaluation framework deeply explores both well-studied and previously unexplored properties, and it uses new metrics in a new comparative experimental setting. By doing so, it adds both depth and width to the current knowledge about the quality of semantic relation knowledge that can be acquired by today’s PLMs.
@@ 2.1 Probing for Hypernymy
In contrast, Ettinger [20] pioneered the study of hypernymy with prompt-based probing. In prompt-based probing, responses from a model are elicited using a prompt, a textual string with slots, all of which are unfilled. An example prompt for hypernymy is “a [w] is a [v]” where [w] is a slot for the word given (called the target word here), whereas [v] will be predicted by a model. As the second word, which is called relatum, depends on the target word and the relation described. A probe is a prompt where the target word has been filled in. For example, Ettinger used the string “a robin is a [v]” as a probe. If a PLM has learned hypernymy well, it should be able to predict hypernyms of “robin” for [v], such as “bird”.
@@
Hanna and Mareček [24] further developed Ettinger’s and Ravichander et al’s prediction task, exploiting the fact that a semantic relation can be expressed by multiple prompts. For the prompts, Hanna and Mareček adopted some lexico-syntactic patterns known from previous work  [25] to be effective at retrieving word pairs in hypernymy and hyponymy. Such prompts include “my favorite [w] is a [v]”, “a [w], such as a [v]”, “a [w] is a type of [v]”, “a [w] is a [v]”. Hanna and Mareček dealt with the problem of ambiguous target words in a different way from Ravichander et al [53]. They attached an example sentence from SemCor [33] to each probe, whether the target word was ambiguous or not. The example sentence was chosen in such a way that the target word was in a specific sense, namely the first WordNet sense whose hypernyms include the gold hypernym. In Hanna and Mareček’s experiments, BERT reached an accuracy of 0.48 in the best setting. Contrary to intuition, the accuracy dropped by about 0.05 when the example sentences were used.
Another modification of the original task is the evaluation data Hanna and Mareček used, which came from category norms. Category norms are items that humans judge to be subtypes of a category given to them. For instance, one category norm for “fish” is “tuna”; “trout” and “salmon” are others. Cohen et al [14] were the first to collect category norms. In their experiment, each participant wrote down category norms included in one of 43 categories given to them, as many as came to mind in 30 seconds. Cohen et al’s original category norm collection was later expanded by Battig and Montague [3], who added 13 new categories (for a total of 56 categories and 2,082 category-category norm pairs). Another collection effort by Overschelde et al [49] added 14 more new categories (for a total of 70 categories and 1,983 category-category norm pairs).
@@
Hanna and Mareček applied the hypernym-based part of Battig and Montague’s category norm data to their probing experiments in an obvious way, using category norms as target words and categories as gold relata. They discarded categories and category norms that were tokenized into multiple tokens, ending up with a total of 863 norms paired with 25 categories.
Recently, large PLMs’ hypernymy knowledge has also gained attention. Tseng et al [65] studied the semantic relation knowledge of Pythia [6] and GPT-3 [8], using prompts consist of two parts: the gloss of the target word [w] and the phrase indicating the relation explicitly, such as “the hypernym of [w] is [v]”. These prompts are meta-linguistic, which means that they could be applied to any relation. This allows Tseng et al to also treat verb hypernymy and noun holonymy. Their metric is the path similarity between the predicted word and the correct answer. The path similarity measures how close a pair of words is in the WordNet taxonomy, where 1 indicates that they are identical and 0 indicates no path exists between them. They tested models on 702 prompts for noun hypernym, 583 for verb hypernym, and 164 for noun holonym. The results showed that, for all models, the path similarities for noun hypernymy are generally higher than those for verb hypernymy and noun holonymy. For example, Pythia-12b achieved 0.51 for noun hypernymy but only 0.28 for verb hypernymy and 0.12 for noun holonymy.
In summary, PLMs demonstrate a reasonable ability to predict hypernyms of target words. However, several interesting avenues have been left underexplored. The most obvious one is that there are several other established semantic relations apart from hypernymy. Although some studies have explored semantic relations other than hypernymy, such as antonymy by Hernandez et al [26], Chanin et al [13] and semantic relatedness by dos Santos and Leal [59], they are still few in number. More importantly, they treated each relation individually. However, as we will show, semantic relations never stand alone.
@@ 2.2 Relations beyond Hypernymy
Hypernymy is just one of the semantic relations, although it is an essential one. Other semantic relations have also been a long-standing research topic in psychological, theoretical, and computational linguistics. Extensive prior work shows that the distinction among such semantic relations is non‑trivial for both humans and language models [11, 12, 31, 60, 47, 2, 70].
In psycholinguistics, Chaffin and Clark [11] researched the similarities and differences between several semantic relations, as perceived by humans. The experiment used a semantic sorting task, where participants are instructed to group together 31 word pairs, each representing a particular semantic relation. The relations came from five broad categories: contrast (including antonymy), similars (including synonymy), class inclusion (including hypernymy), part-whole (holonymy), and case relations (such as the agent-instrument relation and the agent-action relations, exemplified by “farmer”/“tractor” and “dog”/“bark”). The results showed that the human subjects were able to distinguish contrast (including antonymy) most easily from the other four relations. Similars (including synonymy) and class inclusion (including hypernymy) formed a second cluster, whereas case relations and part-whole formed a separate cluster each.
The similarities between hypernymy and holonymy have been extensively discussed in the semantic literature [15, 69, 31]. Both Cruse and Winston et al pointed out that hypernymy and holonymy are similar in that they both involve division and inclusion. Word pairs related by hypernymy and holonymy always consist of a word referring to an entity that undergoes division, with the other word referring to the result of that division, whether as a part or a subclass. Joosten considered both relations under the term denotational inclusion. The similarity between the relations becomes even more obvious when collective nouns are involved. For example, we can say that a table is a kind of furniture, and we can also say that it is a part of furniture, expressions typically associated with hypernymy and holonymy [31].
@@
Synonymy and hypernymy are also closely related, as was confirmed during the creation of Hyperlex [67]. Hyperlex is a lexical resource of semantic relations (hypernymy, hyponymy, meronymy, synonymy, antonymy, co-hyponymy [^2] Co-hyponymy is the relation between two concepts that share a hypernym, such as “hawk” and “robin” whose hypernyms are ”bird”.), holding between 2,616 word pairs of nouns or verbs333290 unrelated word pairs also exist in Hyperlex.. The original pairs in Hyperlex were sampled from WordNet [42] and the University of Southern Florida Norms dataset (USF) [46]. Three human checkers were asked to verify whether the semantic relation holds for each pair sampled; the pair was discarded unless two of them agreed that it did. Next, different crowd workers were asked to assign a score to the remaining pairs, indicating the degree to which the pair satisfies hypernymy. Note that for those pairs that were related in a non-hypernymy relation, the crowd workers should assign a low score. However, the humans’ score for synonymy was close to that for hypernymy pairs if the two words were close to each other in the WordNet hierarchy (i.e., separated by at most two levels).
2.3 Symmetry and Prototypicality
When assessing a model’s knowledge about semantic relations holistically, it is necessary to consider not only if the model uses the semantic relations correctly, but also to what extent it learns what specific properties the semantic relations have.
Symmetry
One such property is symmetry, which is defined as follows: If a word pair $(w_{1},w_{2})$ is in a symmetric relation, then the reverse pair $(w_{2},w_{1})$ is also in the relation.
Symmetry is also a property of some factual knowledge relations. For instance, the factual knowledge relation “is a sibling of” is symmetric. Mruthyunjaya et al [45] proposed metrics in order to assess whether models learn properties of such factual relations, including symmetry. They used prompt-based probing and defined the concept of reciprocal elicitation: for any word pair (target word $w$ and relatum $v$) that forms a symmetric relation, the model should respond with the relatum, when given the target word, and also respond with the target word, when given the relatum. For the probe “Bart Simpson is a sibling of [v]”, they expected models to predict “Lisa Simpson”, and for the converse probe “Lisa Simpson is a sibling of [v]”, to predict “Bart Simpson”.
[…]
@@ 2.3 Symmetry and Prototypicality
where $\mathbb{I}[P]$ is the indicator function that becomes one when $P$ is true and zero otherwise. This score ranges between zero and one.
Intuitively, the symmetry score can be interpreted as the probability of reciprocal elicitation for a semantic relation $r$. A high score means that agents were able to detect symmetry. Mruthyunjaya et al’s results showed that, for symmetry, BERT outperforms even GPT-3 [8].
@@
There have been theoretical discussions of the prototypicality of holonymy. Taylor [63] and Tversky [66] presented top-down accounts, which emphasize that the whole (holonym) is intrinsic in the conceptualization of the part (meronym). They therefore predict that the relationship between the whole and its mandatory parts should be tighter than between the whole and its optional parts. A building is a structure with walls, and walls are defined by their function within a building. This makes “building” a prototypical holonym of “wall”. But not all holonymy pairs are well described by these accounts, because some optional parts also play an important role. This is acknowledged in the bottom-up accounts [34, 41, both cited by Joosten [31]], who state that a whole is formed by assembling a number of other individual wholes, each of which has a separate existence outside the holonymy relation. For example, “sky” is a typical holonym of “cloud” (indeed, in our forthcoming experiments, it happens to be the holonym most frequently mentioned by humans). However, in sunny weather, the sky can be cloudless, so clouds are not intrinsic to the sky. This seems to make bottom-up theories descriptively more adequate, although neither proposes a prototype prediction mechanism.
@@ 3 Methodological Considerations
The majority of the previous research studied hypernymy. Beyond hypernymy, there is also much theoretical and experimental knowledge about semantic relations. Additionally, the literature has established several facts about meta-relations. However, when it comes to practical investigations of model behaviour, it is always only hypernymy that has been studied, even though it is merely one semantic relation among many.
Additionally, we have seen that the relevant theoretical literature has extensively studied meta-relations such as distinguishability between semantic relations (cf. section 2.2). In stark contrast with this, the methodology previously used is unable to establish the degree to which a model mistakes one semantic relation for another, and methodologies for other meta-relations are non-existent. This leaves us with an incomplete understanding of the nature of semantic relations, and of the knowledge that PLMs have about semantic relations.
Therefore, we design new metrics for measuring prototypicality (a property of relations) and distinguishability (a meta-relation). Using these new metrics, and the established ones for symmetry, we study hyponymy, holonymy, meronymy, antonymy and synonymy, as well as hypernymy. We further provide a direct comparison between models and humans on the same task. Such a human ceiling will allow us to interpret the performance of models more meaningfully.
@@
Another recurrent problem for all probing experiments is that most target words from any source, are naturally sense-ambiguous words. Previous experimental attempts to deal with sense ambiguity are suboptimal. Ravichander et al limited the target words they use to the unambiguous words. This limits the research focus to an artificial subset of all possible words and relations, and has the practical disadvantage of considerably reducing the number of target words one can use. The other existing solution is to provide a context of the target word sense, for example in the form of example sentences, as Hanna and Mareček did, but this empirically harmed the performance of models. Our methodology offers an alternative solution to this problem.
@@
Some previous experiments on relatum prediction assumed that there is only one correct relatum for each tuple. Hanna and Mareček [24] acknowledge that “orange” has two gold relata: “color” and “fruit”. They, therefore, defined two separate gold tuples for this probe: (“orange”, HYP, “color”) and (“orange”, HYP, “fruit”). However, when calculating the accuracy scores, they consider only the first item in the responses for both tuples, and then average over the two tuples. In this setting, it is theoretically impossible for a model to achieve the full score (1) for any sense-ambiguous probe, even if the model had the ability to predict both relata. Whether the model predicts [“color”, “fruit”] or [“fruit”, “color”], the accuracy score is always 0.5555In the general case, the highest achievable accuracy score for target words with $n$ relata is $\frac{1}{n}$..
@@
A confounder in the interpretation of numerical results is the use of definite and indefinite determiners in the probes. Previous researchers routinely used probes that include indefinite determiners, such as Ettinger’s (2020) “a robin is a [v]”. The English indefinite determiner changes its form from “a” to “an” if the following word’s pronunciation starts with a vowel. Choosing “a” or “an” in a probe before the [v] slot would therefore bias the prediction by models towards relata with an initial vowel or consonant.
Ettinger ran experiments using both types of determiners, comparing pairs of probes differing only in the determiners used. Manual inspection of the responses showed that BERT indeed always adhered to the morphophonetic rule. For example, given the probe “a hammer is a [v]”, BERT-large predicts [“hammer”, “tool”, “weapon”, “nail”, “device”], whereas given the probe “a hammer is an [v]”, it predicts [“object”, “instrument”, “axe”, “implement”, “explosive”]. This suggests that the BERT models were able to utilize the grammatical information contained in the determiner as a clue.
Ravichander et al, when faced the problem of which indefinite determiner to place before [v], chose to always use the determiner that morphophonetically fits with the gold standard answer. For example, the determiner in the probe “a moth is an [v]” was chosen to be “an” exactly because the gold standard answer was “insect”. However, when the determiner acts as a clue for the model, it becomes impossible to disentangle how much of the results is due to the model’s semantic knowledge and how much to the clue. Choosing randomly also does not solve the problem. Our solution uses both probes with “a” and probes with “an” and merges the results in a statistical manner.
@@
Another understudied aspect is the comparison between different types of PLM models on the semantic probing task. BERT, one of the models widely studied so far, is a masked language model (MLM). MLMs are pretrained on the masked language modeling task, in which a model is asked to recover tokens in a given sentence that are randomly masked. But recently, causal language models (CLMs) such as Llama [64, 18] have shown high performance in many tasks and thus gained attention. CLMs are pretrained on next-token prediction, the task of predicting the rightmost word given a sequence of words. The difference in pretraining tasks means that the models make their decisions based on different kinds of information. MLMs consider the context of both sides of the masked word, while CLMs consider only the preceding context. Previous studies in a number of tasks found a large influence of the type of context used in PLM pretraining on performance. For instance, for factual knowledge, MLMs are superior over CLMs [52, 10, 45]. We are the first to study the role of bidirectional contexts in the recognition of semantic relations. To be fair to both MLMs and CLMs, we have designed all our prompts in such a way that they end with the slot [v].
@@
We always first establish human performance for each task and then use it as the measuring stick for models’ performance.